As a college graduation gift, your Aunt plans to buy shares of two stocks for you. The first costs $23 and she pays $1.40 in dividends per share. The second costs $32 and pays $1.20 per share in dividends. If she has $10,100 to spend and wants you to earn $540 individends, how many of each stock should she buy?
Found 2 solutions by mananth, jorel1380:
Answer by mananth(16946)

$23 stock x numbers 1.4
$32 stock y numbers 1.2

23 x + 32 y = 10100 .............1
1.4 x + 1.2 y = 540 .............2

multiply (1)by -1.2
Multiply (2) by 32
-27.6 x -38.4 y = -12120
44.8 x + 38.4 y = 17280
Add the two equations
17.2 x = 5160
/ 17.2
x = 300
plug value of x in (1)
23 x + 32 y = 10100
6900 + 32 y = 10100
32 y = 10100 -6900
32 y = 3200
y = 100
$23 stock 300 numbers
$32 stock 100 numbers
m.ananth@hotmail.ca

Answer by jorel1380(3719)

23p+32q=10100
140p+120q=54000
345p+480q=151500
560p+480q=216000
215p=64500
p=300
q=100
She should buy 300 shares of the first stock, and 100 shares of the second stock..